Examine the IR spectrum below and identity as molecule (1-chlorohentanel and consider all expe ational modes even in the fing i am not sure how to answer the bottom question...
Answer #1

In the question it is asked that which class of compounds are absent here. We need to exclude them during predicting their names/structures.

Now first need to take a look which class of compound is this if the name of the compound is not given. The peaks obtained around 2800 -2950 cm-1 are the for the C-H stretching frequency of alkanes. The peak came in between 1450-1500 cm-1 (although it marked as 1600 cm-1 in the image) are generally due to CH2 scissoring. We can also see that in the the finger print region there is a small peak which could be characteristics of C-X (X = halogen) stretching, although the peaks in the finger print region is not very reliable. From this data we can qualitatively say that the compound may be an alkyl halide if the name is not given.

Now there are lots of other compounds which have different structures, different functional groups. In the FTIR spectra they also show absorption peaks in different regions based on their properties. Now we need to find out here what types of elements, groups or functional groups are absent here. We have to eliminate those class of compounds from consideration while naming the compound.

For example, alcoholic O-H, alkyne C-H stretching, amine or amide N-H stretching shows peaks in between 3700 cm-1 and 3200 cm-1. Unsaturated C-H (aryl/vinyl) characteristic stretching frequencies comes in between 3100 and 300 cm-1. Alkyne C-C stretching and C-N stretching peaks come in between 2300 - 2000 cm-1. All the carbonyl C=O stretching frequencies of aldehydes, ketones, esters, carboxylic acids, amides come in the region 1750 - 1650 cm-1. Alkene C=C stretching frequency shows appearence generally in the region 1680-1600 cm-1. All of these bands are absent here totally.

So the characteristic peaks of alcohols, alkenes, alkynes, amines, amides, aromatic rings, nitriles, aldehydes, ketones, carboxylic acid, esters, amides are absent here. Therefore all these class of compounds could be eliminated by noting regions where the peaks are absent while predicting the structure/characteristics of the given compound incase the name of the compound was not given.

This is what asked in the question.